OpenStudy (sleepyjess):

What is the rate of change for f(x) = −2 cos 4x − 3 on the interval from x=pi/4 to x=pi/2?

OpenStudy (sleepyjess):

@undeadknight26 can you help?

undeadknight26 (undeadknight26):

o.o this stuff is some really advanced stuff i have not learned yet...sorry mate! @whpalmer4 @iPwnBunnies @surjithayer @hba @Lena772 @Compassionate @mathslover

OpenStudy (sleepyjess):

ok, thank you

OpenStudy (amistre64):

what is the slope of the line between the endpoints of the stated interval?

OpenStudy (amistre64):

for some interval [a,b], the slope between the endpoints is just: \[\frac{f(b)-f(a)}{b-a}\]

OpenStudy (sleepyjess):

That is all the information we were given.

OpenStudy (amistre64):

then use it :)

OpenStudy (amistre64):

they give you a,b, and f(x) .... thats all we need right?

OpenStudy (sleepyjess):

I am so confused with this stuff. I'm in Algebra 2, which I was doing pretty good with, but then the last chapter is trig and I just don't get it.

OpenStudy (amistre64):

do you know how to input a value into a function?

OpenStudy (sleepyjess):

Do I have to graph it first?

OpenStudy (sleepyjess):

yes

OpenStudy (amistre64):

'have to'? no .... it you want to sure ... but all we need is values

OpenStudy (amistre64):

then they gave you f(x), what is f(pi/4)? what is f(pi/2)? what is (pi/4 - pi/2)?

OpenStudy (amistre64):

f(x) = -2cos(4x) - 3 f(pi/4) = -2cos(4pi/4) - 3 f(pi/4) = -2cos(pi) - 3 f(pi/4) = -2(-1) - 3 ... --------------------------- f(x) = -2cos(4x) - 3 f(pi/2) = -2cos(4pi/2) - 3 f(pi/2) = -2cos(2pi) - 3 f(pi/2) = -2(1) - 3

OpenStudy (amistre64):

had the interval going the wrong way but its: (pi/2 - pi/4) = pi/4

OpenStudy (sleepyjess):

f(pi/4)=(0,pi/4)

OpenStudy (amistre64):

f(pi/4) is not a point ... especcially not a point where x=0 and y=pi/4, since x=pi/4

OpenStudy (sleepyjess):

f(pi/2)=(0,pi/2)

OpenStudy (sleepyjess):

Those are the y intercepts

OpenStudy (amistre64):

why in the world would you be trying to assess any y intercepts?

OpenStudy (sleepyjess):

I want to find the slope then?

OpenStudy (amistre64):

the y intercept of a function doesnt change based upon the value of x that you use .... the only x value for a y intercept is x=0: f(0) gives the y intercept value

OpenStudy (amistre64):

you want to find the slope between: f(pi/2) and f(pi/4) yes